okay i have to find the standard form of a line that has slope of 3 and passes through the coordinate pair (4,2)
Answer by chitra(359)

The general equation of the line is given by:

y = mx + b

Where m is the slope and (x,y) is the co-ordinate point.

Let us substitute the given slope and the point in the above equation.

y = (3)x + b

Now substituting the point, we have,

2 = (3)(4) + b

2 = 12 + b

2 - 12 = b

b = -10

Now, we substitute the "b" value and the slope to find the equation of the straight line.

y = 3x - 12

is the equation of the required line.

Hence, the solution.